OpenStudy (anonymous):

Factor this. 6xy+30x-10y-9x^2-25

OpenStudy (aum):

6xy+30x-10y-9x^2-25 = 6xy - 10y - 9x^2 + 30x - 25 = 6xy - 10y - 9x^2 + 30x - 25 = 2y(3x - 5) - (9x^2 - 30x + 25) = 2y(3x-5) - (3x-5)^2 = 2y(3x-5) - (3x-5)(3x - 5) = (3x - 5)(2y - 3x + 5)